Find the number of ways to arrange the word BRAZIL
a)if no repetition is allowed
b)the vowels are next to each other

Answer by Edwin McCravy(20055)

BRAZIL

1. Choose the 1st letter any of 6 ways.
2. Choose the 2nd letter any of the 5 remaining ways. 
That's 6×5 or 30 ways to choose the first 2 letters.
3. Choose the 3rd letter any of the 4 remaining ways. 
That's 6×5×4 or 120 ways to choose the first 3 letters.
4. Choose the 4th letter any of the 3 remaining ways. 
That's 6×5×4×3 or 360 ways to choose the first 4 letters.
5. Choose the 5th letter either of the 2 remaining ways. 
That's 6×5×4×3×2 or 720 ways to choose the first 5 letters.
6. Choose the 6th letter the only 1 remaining way. 
That's 6×5×4×3×2×1 or 720 ways to choose the 6 letters. 

That's 6! = 6×5×4×3×2×1 = 720 ways

Then there are these two cases:

Case 1. Those with AI  (A on the left, I next to A on the right)
Case 2. Those with IA  (I on the left, A next to I on the right) 

Case 1. There are 5 things to arrange B, R, Z, L and AI   

1. Choose the 1st thing any of 5 ways.
2. Choose the 2nd thing any of the 4 remaining ways. 
That's 5×4 or 20 ways to choose the first 2 things.
3. Choose the 3rd thing any of the 3 remaining ways. 
That's 5×4×3 or 60 ways to choose the first 3 things.
4. Choose the 4th thing either of the 2 remaining ways. 
That's 5×4×3×2 or 120 ways to choose the first 4 things.
5. Choose the 5th thing the only 1 remaining way. 
That's 5×4×3×2×1 or 120 ways to choose the 5 things. 

That's 5! = 5×4×3×2×1 = 120 ways

Case 2. There are 5 things to arrange B, R, Z, L and IA

This is the same as case 1, or 120.

Total number of ways = 120+120 = 240 ways.

Edwin